OpenStudy (anonymous):

Simplify.. need help asap... 8/ab + 6/b and (3t2/(t + 2)) x ((t + 2)/t2) thanks!!

jimthompson5910 (jim_thompson5910):

The first one is \[\Large \frac{8}{ab} + \frac{6}{b}\] right?

OpenStudy (anonymous):

yes.

jimthompson5910 (jim_thompson5910):

what is the LCD in this case?

OpenStudy (anonymous):

i don't know.

jimthompson5910 (jim_thompson5910):

we have ab as the first denominator and b as the second denominator

jimthompson5910 (jim_thompson5910):

do you see how the LCD would be ab?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

:( I'm really bad at math. I'm sorry..

jimthompson5910 (jim_thompson5910):

maybe replace a and b with numbers

jimthompson5910 (jim_thompson5910):

say a = 2 b = 3

OpenStudy (anonymous):

okay.

jimthompson5910 (jim_thompson5910):

ab = 2*3 = 6

OpenStudy (anonymous):

i understand that.

jimthompson5910 (jim_thompson5910):

if we had 1/6 + 1/3, what would the LCD be?

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

yes because we can multiply the denominator 3 by 2 to get 6

jimthompson5910 (jim_thompson5910):

we can multiply top and bottom of 1/3 by 2 to get 2/6 1/6 + 1/3 turns into 1/6 + 2/6

jimthompson5910 (jim_thompson5910):

from there you add straight across and leave the denominator alone

OpenStudy (anonymous):

mhm.

jimthompson5910 (jim_thompson5910):

the same idea applies here we multiply top and bottom of the second fraction by 'a' \[\Large \frac{8}{ab} + \frac{6}{b}\] \[\Large \frac{8}{ab} + \frac{6{\color{red}{a}}}{b{\color{red}{a}}}\] \[\Large \frac{8}{ab} + \frac{6a}{ab}\]

jimthompson5910 (jim_thompson5910):

then you add the numerators and place that over the denominator

OpenStudy (anonymous):

okat.

OpenStudy (anonymous):

now I'm lost.

jimthompson5910 (jim_thompson5910):

where at?

OpenStudy (anonymous):

how do we find the answer? or did we already find it and i was lost before then? i thought I've been following.

jimthompson5910 (jim_thompson5910):

do you see why I multiplied top and bottom of the second fraction by 'a'?

jimthompson5910 (jim_thompson5910):

I wanted to get every denominator equal to the LCD 'ab'

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

there's one more step to go

OpenStudy (anonymous):

alright.

jimthompson5910 (jim_thompson5910):

What does \[\Large \frac{8}{ab} + \frac{6a}{ab}\] simplify to?

OpenStudy (anonymous):

8/ab + 6/ab

OpenStudy (anonymous):

i really am lost.

OpenStudy (anonymous):

8/ab+6/b

jimthompson5910 (jim_thompson5910):

just add up the numerators \[\Large \frac{8}{ab} + \frac{6a}{ab}=\frac{8+6a}{ab}\] this is possible because the denominators are both equal to ab

OpenStudy (anonymous):

okay. now how would we do the second one?

jimthompson5910 (jim_thompson5910):

The second one is \[\Large \left(\frac{3t^2}{t+2}\right)\times\left(\frac{t+2}{t^2}\right)\] right?

OpenStudy (anonymous):

t+2 on the sec on one are in a parenthesis.

OpenStudy (anonymous):

and in the first one. its like parenthesis insides of themselves.

jimthompson5910 (jim_thompson5910):

ok so this? \[\Large \left(\frac{3t^2}{(t+2)}\right)\times\left(\frac{(t+2)}{t^2}\right)\]

OpenStudy (anonymous):

yes that is correct.

jimthompson5910 (jim_thompson5910):

first notice this cancellation \[\Large \left(\frac{3{\color{red}{\cancel{\color{black}{t^2}}}}}{t+2}\right)\times\left(\frac{t+2}{{\color{red}{\cancel{\color{black}{t^2}}}}}\right)\] the t^2 terms will go away leaving just \[\Large \left(\frac{3}{(t+2)}\right)\times\left(\frac{(t+2)}{1}\right)\]

OpenStudy (anonymous):

ok what is the next step

jimthompson5910 (jim_thompson5910):

then the (t+2) terms also cancel \[\Large \left(\frac{3}{{\color{red}{\cancel{\color{black}{(t+2)}}}}}\right)\times\left(\frac{{\color{red}{\cancel{\color{black}{(t+2)}}}}}{1}\right)\]

jimthompson5910 (jim_thompson5910):

I'm sure you see how to finish

OpenStudy (anonymous):

so its 3?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thats the easiest thing I've ever seen